OpenStudy (lataesiaaa):

i need help to see which one of these inverse functions match. ill post them in the komments

OpenStudy (lataesiaaa):

f^-1(x)=2x-3/5 .

OpenStudy (emyoung0):

what are the other options?

OpenStudy (lataesiaaa):

1) f(x)=2x-8 2) f(x)=2x+7 3) f(x)=7x-2 4) f(x)=2x+8/3 5) f(x)=5x+3/2 6) f(x)=1/x+2